Chapter 24, Probie 3/15/2018 Print by: MELISSA SIMION Physics 202-001 Spring 2018/ DiFabio Phys 202 Chap 24 part 2 Chapter 24, Problem 041 e point x 62 cm on an x axis. The particle A particle of charge +3.4 μC is released from rest at th begins to move due to the presence o kinetic energy of the p HC? f a charge Q that remains fixed at the origin. What is the article at the instant it has moved 16 cm if (a)Q = +55 μC and (b)Q =-55 (a) Number *1 (b) Number Units *2 Answer *1: the tolerance is +/-200 Answer * 2: the tolerance is +/-2% Question Attempts: Unlimited
Answer #1

here,

charge , Q1 = 3.4 * 10^-6 C is at x1 = 62 cm = 0.62 m

a)

when Q = 55 * 10^-6 C

r = 0.16 m

let the kinetic energy be KE

KE = K * Q * Q1 * ( 1 /x1 - 1/(x1 + r))

KE = 9 * 10^9 * 3.4 * 10^-6 * 55 * 10^-6 * ( 1/0.62 - 1/0.78)

KE = 0.56 J

b)

when Q = - 55 * 10^-6 C

r = 0.16 m

let the kinetic energy be KE

KE = K * Q * Q1 * ( 1 /(x1 - r ) - 1/(x1))

KE = 9 * 10^9 * 3.4 * 10^-6 * 55 * 10^-6 * ( 1/(0.62 - 0.16) - 1/0.62)

KE = 0.94 J